# Implementation of full and simplified likelihoods in CheckMATE Iñaki Lara and Krzysztof Rolbiecki

### v2

@@ 1 Introduction
The Run-3 at the LHC is now in full swing and by the end of the current run it is expected that the LHC would collect several hundreds of inverse femtobarns of data. Apart from the Higgs boson discovery[1, 2], which made the experimental picture of the Standard Model complete, no sign of new physics has been observed. However, the vast data has only been interpreted in the limited number of theoretical models by the LHC experiments. On many occasions, the interpretation is provided in terms of simplified models [3] which do not directly correspond to realistic TeV-scale physics. This situation has prompted several groups [4, 5, 6, 7, 8, 9] to provide computer programs that would allow reinterpretation of experimental results in terms of an arbitrary new physics model [10].
Over the years, searches for new physics were becoming increasingly sophisticated, including the full use of complicated statistical models [11] and machine learning methods [12, 13]. Instead of reporting results in a handful of bins, often optimized for a small class of models, and comparing observed and expected number of events, the experiments nowadays provide binned data in several observables in signal and control regions. On the one hand, this can significantly improve the sensitivity, but on the other hand, it greatly complicates the reinterpretation of the results. In this approach, experiments in high-energy physics test the compatibility of collision data with theoretical predictions, which can be described as a likelihood function. The function gives the probability of the data assuming a theoretical model and a certain set of parameters. The likelihood can therefore be used to constrain the parameters of the model.
In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [14, 15, 16, 17]. For implementations in other tools, see eg. [18, 19]. In this paper, we include 9 ATLAS and 4 CMS searches. The CMS searches use the correlated background model (covariance matrix) [20]. For ATLAS searches, we use the simplified [21] and the full likelihood [22] frameworks. In addition, some of the implementations of the ATLAS searches include the full set of control regions. Users have several ways of controlling the actual calculation of likelihoods. The default method of evaluation is based on the package Spey [23]. We provide validation plots for each of the searches. A comparison of three methods for the calculation of limits is presented: the best-expected signal region, the full likelihood model (if available), and the simplified likelihood.
@@ 2.1 ATLAS
The functionality of combining signal regions for recasting in ATLAS searches can be implemented using either the full likelihood model [22] or following a simplified approach detailed in Ref. [21]. Table 1 lists the ATLAS analyses with likelihood functionality implemented in CheckMATE. The simplified likelihood method requires background rates and uncertainties that were already available in the implemented searches. The full likelihood requires an appropriate file in the JSON format and these files were released by ATLAS for 6 searches already implemented in CheckMATE. For the searches atlas_2004_14060, atlas_2006_05880, and atlas_2111_08372 the full model files are not available but using the published data one can still perform simplified model fitting in multibin signal regions.
@@
The full-likelihood statistical models are encoded in the JSON files by the ATLAS Collaboration. The files are not shipped with CheckMATE but are automatically downloaded during installation from the HEPData website, https://www.hepdata.net/. The information provided includes the number of background events for all signal and control regions and for each major background category separately. This results in a large number of nuisance parameters and the complexity of the procedure makes the hypothesis testing very CPU-expensive. Additionally, on the recasting side, in order to fully exploit the method, one should also implement control regions (CRs), which was not a standard approach in CheckMATE. Currently, just two searches atlas_2010_14293 and atlas_1911_06660 has a complete implementation of all CRs. In other searches, it is assumed that the contribution of signal to CRs is negligible. This assumption is not obviously fulfilled in all imaginable new physics models.
@@
On the technical side, after the usual evaluation of events within CheckMATE, a JSON patchset is created that encapsulates signal contributions to signal regions (SRs) (and CRs if applicable). The patchset is then combined with the background-only input from ATLAS. Since the signal region names in CheckMATE do not usually exactly match the names in ATLAS conventions, an additional file is required, pyhf_conf.json, which provides a dictionary between two conventions. This is further evaluated using the package Pyhf [24, 25, 26], which is a Python implementation of the HistFactory specification for binned statistical models [27, 28]. The signal strength $\mu$ is the parameter of interest, where $\mu=1$ corresponds to the nominal cross section of a tested model. Depending on the user’s choices, the output can contain information about the expected and observed upper limit on $\mu$, with 2-$\sigma$ bounds, along with the observed and expected $\mathrm{CL}_{s}$ for $\mu=1$. The default method of calculation is by using the asymptotic formulae; see [24].
[…]
Since the evaluation of full likelihoods is normally time-consuming, it is not practical for large scans of the parameter space. Therefore, the alternative approach to likelihood evaluation is based on the concept of simplified likelihood [21]. In this case, the background model is approximated with the total SM background rate obtained in the background-only fit in the full model. A single nuisance parameter correlated over all bins and representing post-fit background uncertainty is constrained by the unit normal distribution. The evaluation is also performed using the Pyhf package.
Name Description #SR Nbin Full Ref.
@@
[…]
atlas_2111_08372 Search for associated production of a Z boson with an invisibly decaying Higgs boson or dark matter candidates $2 \usym1$F5F4 [39]
@@ 2.2 CMS
where the product runs over all bins and $\mu$ is the signal strength (and the Parameter of Interest - POI), $n_{i}$ the observed number of events, $s_{i}$ an expected number of signal events, $b_{i}$ an expected number of background events, $\theta_{i}$ a background nuisance parameter and $\mathbf{V}$ the covariance matrix. It is implemented using the covariance matrices provided by the CMS Collaboration which are included in the CheckMATE distribution in the JSON format. Evaluation of the above model is performed using the Spey package and the default_pdf.correlated_background method.
@@ 2.3 CheckMATE parameters
CheckMATE provides several switches and parameters to control details of statistical evaluation. These are summarized in Tab. 3. The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation. By default, no statistical evaluation is performed. For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information. Generally, the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures. By choosing a select switch users can control which statistics are calculated. If no explicit choice is made, the observed upper limit will be calculated. Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming. The option -so can be used to request calculation of statistics for previous CheckMATE runs (it requires presence of the evaluation/total_results.dat file in the output directory).
The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section). For the default method, simple, calculation is performed using simplified likelihood and the CheckMATE-Pyhf interface. The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface. Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible regarding the output compared to the previous options). In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
Model: X -mod X simple The simplified likelihood model for ATLAS searches (default).
full The Speyinterface to the full likelihood model for ATLAS searches.
@@
The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file. In order to follow the progress of the calculation, the observed limits are also displayed on screen for each of the signal regions. The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
$\mu<1\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
@@
If the results for upper limit are not available the decision is made using the observed $\mathrm{CL}_{s}$ statistics at the 95% confidence level:
$\mathrm{CL}_{s}<0.05\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
@@ 2.4 Evaluation time
Four different backends for calculation of full likelihood models in the pyhf/Spey setup are supported: Numpy [43], Tensorflow [44], Pytorch [45], and Jax [46]. The Numpy backend does not require additional system components to be installed but it generally is not recommended due to long computation time. When no backend is explicitly selected, CheckMATE will check if the Jax backend is available. The Jax backend was also successfully tested with NVIDIA CUDA library and GPU support.
@@
As already mentioned, for large scans of parameter space it may not be practical to perform full likelihood evaluation. The simplified models improve computation and the detailed calculation of all signal strength limits and $\mathrm{CL}_{s}$ takes in our example 16 seconds for the Jax backend and 64 seconds for the Numpy backend. Finally, if the calculation of $\mathrm{CL}_{s}$ is sufficient, it only takes about 1 second.
@@ 3 Validation
Multibin signal regions are currently available in 9 ATLAS and 4 CMS analyses, as listed in Tabs. 1 and 2. The searches are based on the full Run 2 luminosity of about 140 ${\text{fb}}^{-1}$ at the center-of-mass energy $\sqrt{s}=13\ \mathrm{TeV}$. In this section we briefly introduce each of the searches and provide validation examples. When possible we compare full vs. simplified likelihood approach, and provide examples how the multibin approach improves exclusions compared to the best-signal-region approach. The best-SR (BSR) is defined as the signal region with the best expected sensitivity.
@@
The validation procedure is organized as follows. For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state. The NNPDF23LO [50, 51, 52] parton distribution function (PDF) set is used. The events are then interfaced to Pythia 8.3 [53, 54] for modeling of decays, hadronization and showering. The matrix element and parton shower matching for ATLAS searches was performed using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set. For CMS searches, the MLM [56, 48] matching scheme was used. Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], following recommendations of Ref. [68]. The signal cross sections for production of sleptons, charginos and neutralinos are computed at next-to-leading order plus next-to-leading-log precision using Resummino [69, 70, 71, 72, 73, 74, 75]. This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
@@
Events for the simplified scenario of DM production in the CMS search EXO-20-004 [42] were generated at NLO with MadGraph5_aMC@NLO 3.1.0 and the DMsimp model [76]. Up to 2 additional partons were included and the FxFx [77] jet matching scheme was used. Samples for the 2HDM+a model and validation of HIGG-2018-26 [39] were simulated internally in CheckMATE interfaced to MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
@@
For the purpose of validation, we present comparisons between CheckMATE and official ATLAS result of both expected and observed exclusion limits. Where available, the results for simplified and full likelihood are provided. In addition, we also include exclusion contours using the best expected signal region. Note that this usually is based on comparing observed signal with the post-fit background-only prediction in a given bin, which typically has reduced uncertainty, and hence can be quite an aggressive approach. Also, a limit obtained with the full likelihood model is not a combination of such bin-by-bin measurements and cannot be compared in a straightforward way. On the other hand, the simplified likelihood will be such a combination in most cases and it is not surprising that it will produce stronger constraints than the full likelihood in some searches, see also Ref. [23] The latter generally shows weaker sensitivity than multibin histogram shape-fits, but can be advantageous in terms of evaluation speed.
@@
We follow the same key for the plots throughout this section. The black solid (dashed) line is for ATLAS/CMS observed (expected) limit. On some occasions, this is accompanied by theoretical (experimental) $\pm 1\sigma$ uncertainty. The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively. Finally, the blue and green lines denote the limits obtained using the simplified and full likelihoods, respectively.
@@
This is a search [31] for bottom squark production in final states containing Higgs bosons, $b$-jets, and missing transverse momentum. The Higgs boson is reconstructed from two $b$-tagged jets. The final states contain at least 3 (SRC) or 4 (SRA, SRB) $b$-jets, no leptons, and large missing transverse momentum. The signal region A is divided into 3 bins according to effective mass, $m_{\textrm{eff}}$, and the signal region C is divided into 4 bins of missing transverse energy significance, $\mathcal{S}$. Thus, both SRA and SRC allow for a shape-fit analysis. The full likelihood model is provided.
The validation plots for this search are presented in Fig. 1. We compare bottom squark pair production, $pp\to\tilde{b}_{1}\tilde{b}_{1}^{*}$, and two distinct mass spectra. The first assumes that the mass of the lightest supersymmetric particle (LSP) is $m_{\tilde{\chi}_{1}^{0}}=60$ GeV and the other assumes the mass difference between neutralinos $\Delta m(\tilde{\chi}_{2}^{0},\tilde{\chi}_{1}^{0})=130$ GeV. Generally a good agreement between CheckMATE and ATLAS is observed for the shape-fit method with CheckMATE being slightly weaker. Notably, there is a very good agreement between simplified and full models. The best SR method performs well for the model with fixed LSP mass, however, in the second scenario its exclusion strength can be up to 3 times weaker than shape-fits.
@@
This is a search [32] for production of electroweakinos and sleptons. The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered. Events are first separated into SF and DF categories and are further subdivided by the multiplicity of the non-b-tagged jets. A combined multi-bin SR is defined out of the 36 exclusive binned signal regions. The full likelihood statistical model is provided.
The validation plots for the production of slepton pairs are shown in Fig. 2. Both the full- and simplified likelihood models agree very well with the ATLAS result. The BSR method gives visibly weaker constraints.
@@
This is a search [33] for production of staus in final states with two hadronic $\tau$-leptons and missing transverse momentum. Two orthogonal signal regions can be combined in a fit for which the full likelihood model is provided. The CheckMATE implementation includes one of the multijet control regions (CR-A), for which a significant contribution is expected, up to 30%, from signal events.
The validation plots in Fig. 3 show a similar sensitivity of the best signal region, as well as full and simplified likelihoods. In each case, good agreement with the ATLAS result is observed. The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably within the experimental $1\sigma$ uncertainty. It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
Figure 3: Validation plots for the search atlas_1911_06660 (SUSY-2018-04). Left panels: observed limits; right panels: expected limits.

### v3

# ul. Pasteura 5, PL-02-093 Warsaw, Poland

@@ 1 Introduction
The Run 3 at the Large Hadron Collider (LHC) is now in full swing, and by the end of the current run it is expected that the LHC would collect several hundreds of inverse femtobarns of data. Apart from the Higgs boson discovery[2, 52], which made the experimental picture of the Standard Model complete, no sign of new physics has been observed. However, the vast data has only been interpreted in the limited number of theoretical models by the LHC experiments. On many occasions, the interpretation is provided in terms of simplified models [20] which do not directly correspond to realistic TeV-scale physics. This situation has prompted several groups [78, 17, 56, 43, 30, 99] to provide computer programs that would allow reinterpretation of experimental results in terms of an arbitrary new physics model [13].
Over the years, searches for new physics were becoming increasingly sophisticated, including the full use of complicated statistical models [59] and machine learning methods [15, 25]. Instead of reporting results in a handful of bins, often optimized for a small class of models, and comparing observed and expected numbers of events, the experiments nowadays provide binned data in several observables in signal and control regions. On the one hand, this can significantly improve the sensitivity, but on the other hand, it greatly complicates the reinterpretation of the results. In this approach, experiments in high-energy physics test the compatibility of collision data with theoretical predictions, which can be described as a likelihood function. The function gives the probability of the data assuming a theoretical model and a certain set of parameters. The likelihood can therefore be used to constrain the parameters of the model.
In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [65, 77, 63, 64]. For implementations in other tools, see, e.g. [18, 16]. In this paper, we include 9 ATLAS and 4 CMS searches. The CMS searches use the correlated background model (covariance matrix) [55]. For ATLAS searches, we use the simplified [74] and full likelihood [88] frameworks. In addition, some of the implementations of the ATLAS searches include the full set of control regions. Users have several ways of controlling the actual calculation of likelihoods. The default method of evaluation is based on the Spey package [26]. We provide validation plots for each of the searches. A comparison of three methods for the calculation of limits is presented: the best-expected signal region, the full likelihood model (if available), and the simplified likelihood.
@@ 2.1 ATLAS
The functionality of combining signal regions for recasting in ATLAS searches can be implemented using either the full likelihood model [88] or following a simplified approach detailed in Ref. [74]. Table 1 lists the ATLAS analyses with likelihood functionality implemented in CheckMATE. The simplified likelihood method requires background rates and uncertainties that were already available in the implemented searches. The full likelihood requires an appropriate file in the JSON format and these files were released by ATLAS for 6 searches already implemented in CheckMATE. For the searches atlas_2004_14060, atlas_2006_05880, and atlas_2111_08372 the full model files are not available but using the published data one can still perform a simplified model fitting in multibin signal regions.
@@
The full-likelihood statistical models are encoded in the JSON files by the ATLAS Collaboration. The files are not shipped with CheckMATE but are automatically downloaded during installation from the HEPData website [48, 83], https://www.hepdata.net/. The information provided includes the number of background events for all signal and control regions and for each major background category separately. This results in a large number of nuisance parameters and the complexity of the procedure makes the hypothesis testing very CPU-expensive. Additionally, on the recasting side, in order to fully exploit the method, one should also implement control regions (CR), which was not a standard approach in CheckMATE. Currently, only two searches atlas_2010_14293 and atlas_1911_06660 have complete implementation of all CRs. In other searches, it is assumed that the contribution of signal to CRs is negligible. This assumption is not obviously fulfilled in all imaginable new physics models.
@@
On the technical side, after the usual evaluation of events within CheckMATE, a JSON patchset is created that encapsulates signal contributions to signal regions (SRs) (and CRs if applicable). The patchset is then combined with the background-only model from ATLAS. Since the signal region names in CheckMATE do not usually exactly match the names in ATLAS conventions, an additional file is required, pyhf_conf.json, which provides a dictionary between two conventions. The likelihood is further evaluated using the Pyhf [57, 72, 73] package, which is a Python implementation of the HistFactory specification for binned statistical models [58, 28]. The signal strength $\mu$ is the parameter of interest, where $\mu=1$ corresponds to the nominal cross section of a tested model. Depending on the user’s choices, the output can contain information about the expected and observed upper limit on $\mu$, with 2-$\sigma$ bounds, along with the observed and expected $\mathrm{CL}_{s}$ for $\mu=1$, where
$\mathrm{CL}_{s}=\frac{\mathrm{CL}_{s+b}}{\mathrm{CL}_{b}}$ (1)
and $\mathrm{CL}_{s+b}$, $\mathrm{CL}_{b}$ are the $p$-values of the signal and the null hypotheses, respectively. Asymptotic formulae are the default calculation method; see [57].
@@
Name Description #SR Nbin Full Ref.
atlas_1908_03122 Search for bottom squarks in final states with Higgs bosons, $b$-jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 2 7
[3]
atlas_1908_08215 Search for electroweak production of charginos and sleptons in final states with 2 leptons and $E_{\mathrm{T}}^{\mathrm{miss}}$ 1 52
[6]
atlas_1911_06660 Search for direct stau production in events with two hadronic taus 1 2
[5]
atlas_1911_12606 Search for electroweak production of supersymmetric particles with compressed mass spectra 2 76
[8]
atlas_2004_14060 Search for stops in hadronic final states with $E_{\mathrm{T}}^{\mathrm{miss}}$ 3 14
[4]
atlas_2006_05880 Search for top squarks in events with a Higgs or $Z$ boson 3 23
[7]
atlas_2010_14293 Search for squarks and gluinos in final states with jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 3 60
[9]
atlas_2101_01629 Search for squarks and gluinos in final states with one isolated lepton, jets, and $E_{\mathrm{T}}^{\mathrm{miss}}$ 1 26
[10]
atlas_2111_08372 Search for associated production of a $Z$ boson with an invisibly decaying Higgs boson or dark matter candidates 1 22
[11]
@@ 2.2 CMS
where the product runs over all bins and $\mu$ is the signal strength (and the Parameter of Interest - POI), $n_{i}$ the observed number of events, $s_{i}$ the expected number of signal events, $b_{i}$ the expected number of background events, $\theta_{i}$ a background nuisance parameter, and $\mathbf{V}$ the covariance matrix. It is implemented using the covariance matrices provided by the CMS Collaboration, which are included in the CheckMATE distribution in the JSON format. The evaluation of the above model is performed using the Spey package and the default_pdf.correlated_background method.
@@ 2.3 CheckMATE parameters
CheckMATE provides several switches and parameters to control the details of the statistical evaluation. These are summarized in Tab. 3. The switches are divided into two groups: one providing a control of what statistical tests are performed, and the other to control different modes of calculation. By default, no statistical evaluation is performed. For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining an Allowed/Excluded result but with limited additional information. Generally, available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures. By choosing a select switch, users can control which statistics are calculated. If no explicit choice is made, the observed upper limit will be calculated. Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming. The option -so can be used to request the calculation of statistics for previous CheckMATE runs (it requires the presence of the evaluation/total_results.dat file in the output directory).
The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section). For the default method, the full switch chooses the calculation using the full likelihood and the CheckMATE-Spey interface. With the simple switch, the calculation is performed using simplified likelihood and the CheckMATE-Pyhf interface. Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible with regard to the output compared to the previous options). In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
full The Speyinterface to the full likelihood model for ATLAS searches (default).
Model: X -mod X simple The simplified likelihood model for ATLAS searches.
@@
The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file. In order to follow the progress of the calculation, the observed limits are also displayed on the screen for each of the signal regions. The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
$\mu<1\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Excluded}$
@@
If the results for the upper limit are not available, the decision is made using the observed $\mathrm{CL}_{s}$ statistics at the 95% confidence level:
$\mathrm{CL}_{s}<0.05\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Excluded}$
@@ 2.4 Evaluation time
Four different backends for the calculation of full likelihood models in the pyhf/Spey setup are supported: Numpy [86], Tensorflow [96], Pytorch [87], and Jax [76]. The Numpy backend does not require additional system components to be installed, but it generally is not recommended due to long computation time. When no backend is explicitly selected, CheckMATE will check if the Jax backend is available. The Jax backend was also successfully tested with NVIDIA CUDA library and GPU support.
@@
As already mentioned, for large scans of parameter space it may not be practical to perform full likelihood evaluation. The simplified models improve the computation and detailed calculation of all signal strength limits and $\mathrm{CL}_{s}$ takes in our example 16 seconds for the Jax backend and 64 seconds for the Numpy backend. Finally, if the calculation of $\mathrm{CL}_{s}$ is sufficient, it only takes about 1 second.
@@ 3 Validation
Multibin signal regions are currently available in 9 ATLAS and 4 CMS analyses, as listed in Tabs. 1 and 2. The searches are based on the full Run 2 luminosity of about 140 ${\text{fb}}^{-1}$ at the center-of-mass energy $\sqrt{s}=13\ \mathrm{TeV}$. In this section, we briefly introduce each of the searches and provide validation examples. When possible, we compare the full and simplified likelihood approaches, and provide examples of how the multibin evaluation improves exclusion limits compared to the best-signal-region approach. The best-SR (BSR) is defined as the signal region with the best expected sensitivity.
@@
The validation procedure is organized as follows. For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [21, 23, 22] and the MSSM_SLHA2 [66] UFO [60] model with up to two additional partons in the final state. The NNPDF23LO [31, 47, 32] parton distribution function (PDF) set is used. The events are then interfaced with Pythia 8.3 [91, 44] to model decays, hadronization, and showering. The matrix element and parton shower matching for ATLAS searches was performed using the CKKW-L [82] prescription and a matching scale of 1/4 of the SUSY particle mass was set. For CMS searches, the MLM [84, 23] matching scheme was used. In the final step, a fast detector simulation is performed using Delphes 3.5 [61]. Jets are reconstructed using FastJet [51] and the anti-$k_{t}$ algorithm [50].
Inclusive signal cross sections for the production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at the next-to-next-to-leading-logarithm order (approximate NNLO+NNLL) [38, 34, 80, 79, 39, 41, 42, 37, 35, 40, 36], following the recommendations of Ref. [49]. The signal cross sections for production of sleptons, charginos and neutralinos are computed at next-to-leading order plus next-to-leading-log precision using Resummino [71, 68, 46, 45, 67, 70, 62]. This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
@@
Events for the simplified scenario of DM production in the CMS search EXO-20-004 [97] were generated at NLO with MadGraph5_aMC@NLO 3.1.0 and the DMsimp model [29]. Up to 2 additional partons were included and the FxFx [69] jet matching scheme was used. The samples for the 2HDM+a model [33] and the validation of HIGG-2018-26 [11] were simulated internally in CheckMATE interfaced with MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
The sampling of parameter space generally follows grids provided by experiments in the supplementary materials on HEPData. The number of Monte-Carlo (MC) events per point varies; however, close to exclusion lines at mass limits, we aim at samples 10-20 times larger than the nominal luminosity derived values. The exclusion lines are then interpolated based on the grids of simulated parameter points, with additional sampling points occasionally added. MC errors are consistently included in the pyhf patch files.
@@
For the purpose of validation, we present comparisons between CheckMATE and the official ATLAS result of both expected and observed exclusion limits. Where available, the results for simplified and full likelihood are provided. In addition, we also include exclusion contours using the best-expected signal region. Note that this usually is based on comparing observed signal with the post-fit background-only prediction in a given bin, which typically has reduced uncertainty, and hence can be quite an aggressive approach. Also, a limit obtained with the full likelihood model is not a combination of such bin-by-bin measurements and cannot be compared in a straightforward way. On the other hand, the simplified likelihood will be such a combination in most cases and it is not surprising that it will produce stronger constraints than the full likelihood in some searches; see also Ref. [26] The latter generally shows weaker sensitivity than multibin histogram shape-fits, but can be advantageous in terms of evaluation speed.
@@
We follow the same key for the plots throughout this section. The black solid (dashed) line is for the observed (expected) ATLAS/CMS limit. On some occasions, this is accompanied by theoretical (experimental) $\pm 1\sigma$ uncertainty. The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for the observed and expected limits, respectively. Finally, the blue and green lines denote the limits obtained using the simplified and full likelihoods, respectively.
@@
This is a search [3] for bottom squark production in final states containing Higgs bosons, $b$-jets, and missing transverse momentum. The Higgs boson is reconstructed from two $b$-tagged jets. The final states contain at least 3 (SRC) or 4 (SRA, SRB) $b$-jets, no leptons, and large missing transverse momentum. The signal region A is divided into 3 bins according to effective mass, $m_{\textrm{eff}}$, and the signal region C is divided into 4 bins of missing transverse energy significance, $\mathcal{S}$. Thus, both SRA and SRC allow for a shape-fit analysis. Three full likelihood models are provided for the signal regions SRA, SRB, and SRC.
The validation plots for this search are presented in Fig. 1. We compare bottom squark pair production, $pp\to\tilde{b}_{1}\tilde{b}_{1}^{*}$, and two distinct mass spectra. The first assumes that the mass of the lightest supersymmetric particle (LSP) is $m_{\tilde{\chi}_{1}^{0}}=60$ GeV and the other assumes the mass difference between neutralinos $\Delta m(\tilde{\chi}_{2}^{0},\tilde{\chi}_{1}^{0})=130$ GeV. Generally a good agreement between CheckMATE and ATLAS is observed for the shape-fit method with CheckMATE being slightly weaker. In particular, there is a very good agreement between simplified and full models. The best SR method performs well for the model with fixed LSP mass, however, in the second scenario its exclusion strength can be up to 3 times weaker than shape-fits.
@@
This is a search [6] for the production of electroweakinos and sleptons. The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered. Events are first separated into SF and DF categories and are further subdivided by the multiplicity of the non-b-tagged jets. A combined multi-bin SR is defined out of the 36 exclusive binned signal regions. The full likelihood statistical model is provided.
The validation plots for the production of slepton pairs are shown in Fig. 2. Both the full- and simplified likelihood models agree very well with the ATLAS result. The BSR method gives visibly weaker constraints. Additional validation material for CheckMATE, including cut-flows, can be found in [95].
@@
This is a search [5] for the production of staus in final states with two hadronic $\tau$-leptons and missing transverse momentum. Two orthogonal signal regions can be combined in a fit for which the full likelihood model is provided. The implementation includes one of the multijet control regions (CR-A), for which a significant contribution is expected, up to 30%, from signal events.
The validation plots in Fig. 3 show a similar sensitivity of the best signal region, as well as full and simplified likelihoods. In each case, good agreement with the ATLAS result is observed. The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably within the experimental $1\sigma$ uncertainty. A comparison of observed 95% CL upper limits on cross section for several representative parameter points is shown in Tab. 5. In the bottom plot of Fig. 3 we additionally compare the expected limits obtained with and without the inclusion of control regions. Note that for this search, the signal contribution to the control regions is not negligible. Without control regions, the search over-constrains the parameter space, although the discrepancy remains within the 1-$\sigma$ expected uncertainty.
Figure 3: Validation plots for the search atlas_1911_06660 (SUSY-2018-04). Left panels: observed limits; right panels: expected limits. In the bottom row: a comparison between the observed limits derived with and without control regions.
$(m_{\tilde{\tau}_{1}},m_{\tilde{\chi}_{1}^{0}})$ [GeV] full simple ATLAS
$(320,120)$ $3.7$ $4.0$ $4.6$
$(320,160)$ $5.2$ $5.8$ $5.0$
$(400,80)$ $1.9$ $1.6$ $2.0$
$(440,80)$ $1.5$ $1.4$ $1.7$
Table 5: Comparison of observed 95% CL upper limits on cross section in fb for stau pair production in the search atlas_1911_06660 (SUSY-2018-04); full and simple columns correspond to the limits obtained with CheckMATE using the full and simplified likelihood, respectively.